# Deep learning in a bilateral brain with hemispheric specialization

### v8

@@ Abstract
The brains of all bilaterally symmetric animals on Earth are divided into left and right hemispheres. The anatomy and functionality of the hemispheres have a large degree of overlap, but there are asymmetries and they specialise to possess different attributes. Other authors have used computational models to mimic hemispheric asymmetries with a focus on reproducing human data on semantic and visual processing tasks. We took a different approach and aimed to understand how dual hemispheres in a bilateral architecture interact to perform well in a given task. We propose a bilateral artificial neural network that imitates lateralisation observed in nature: that the left hemisphere specialises in specificity and the right in generality. We used different training objectives to achieve the desired specialisation and tested it on an image classification task with two different CNN backbones – ResNet and VGG. Our analysis found that the hemispheres represent complementary features that are exploited by a network head which implements a type of weighted attention. The bilateral architecture outperformed a range of baselines of similar representational capacity that don’t exploit differential specialisation, with the exception of a conventional ensemble of unilateral networks trained on a dual training objective for specifics and generalities. The results demonstrate the efficacy of bilateralism, contribute to the discussion of bilateralism in biological brains and the principle may serves as an inductive bias for new AI systems.
Keywords Bilateral  $\cdot$ Hemispheric specialisation  $\cdot$ Hemispheric asymmetry  $\cdot$ Brain-inspired architecture  $\cdot$ Cognitive architectures  $\cdot$ Deep learning  $\cdot$ Computational cognitive neuroscience  $\cdot$ Inductive bias
@@ 1 Introduction
Left and right hemispheres are a remarkably conserved feature of brains across species, suggesting their importance for intelligence. The anatomy and functionality of the hemispheres have a large degree of overlap, but they are asymmetric and specialise to possess different attributes. For example, the left hemisphere is specialised for specific (fine) classes and familiarity/routine whereas the right hemisphere is specialised for more general (coarse) classes and novelty [4, 5].
Researchers have studied the bilateral brain with computational models that mimic anatomical and functional asymmetries. The models are typically subjected to standard tasks and compared to human behavioural data. The various studies provide evidence that asymmetric functional and anatomical differences can give rise to observed lateralisation of activity (i.e. specialisation) for specific tasks e.g. [1, 18, 13, 8, 11].
The central question of existing models is to understand the cause of the observed specialisation. The aim of this project is to investigate a different question – what is the benefit of specialised hemispheres and how do they complement each other for a given task? Our approach is to induce specialisation directly, rather than test if it emerges as a result of architectural asymmetry. Our objective is to develop a theory of operation and to understand how bilateral principles could be exploited for AI/ML.
1.1 Our approach
We focused on the left hemisphere’s specialisation on specifics, and the right’s specialisation on generalities [4, 5]. We hypothesised that an ANN with differentially specialised sub-networks would outperform a single comparable network in a classification task featuring both fine and coarse class labels. We predicted that with specialisation, the hemispheres would extract distinct but overlapping features, despite identical observations. In addition, we expected that a bilateral architecture would be able to combine these features to improve overall accuracy.
We chose a hierarchical image classification task where each sample belongs to both a fine and a coarse class e.g. coarse: sea creature, fine: penguin, seal, shark. We modelled the hemispheres with left and right CNNs and induced specialisation using supervised training with different objectives, using either coarse or fine labels. The bilateral network was compared to several comparative baselines that do not have specialisation but do have the same number of learnable parameters. The differences were analysed to explore the effects of bilateral specialisation. Experiments were repeated with two types of CNN hemispheres, VGG and ResNet variants, to test generalisation over different backbones.
@@ 2 Related work
Biologically inspired multi-network architectures
Beaulieu et al. [20] created a dual-stream architecture called Neuromodulator Meta-learner, where one network learns to modulate the other, to enhance continual learning. Bakhtiari et al. [21] created a network with parallel pathways to reproduce the functionality of dorsal (‘where’) and ventral (‘what’) pathways in an ANN trained with a single loss function. Li and Deza [22] discussed specialisation of a branched neural network.
In the most related study, Mayan et al. [19] tackled the same image classification task with fine and coarse (hierarchical) classes. The authors created a single network with two hemispheres, and trained it with supervised learning and a single loss function. They used hyperparameters with analogies to biological parameters in each hemisphere to encourage specialisation. Specialisation was achieved in each side but the bilateral network did not have an advantage over baselines.
Ensemble models
@@
Multi-domain learning
In multi-domain learning (MDL), multiple models are trained to specialise in a particular domain [27]. Two of the main approaches include a) modelling domain-specific and domain-general features [28] and b) explicitly modelling the relationship between domains. MDL is focused on classes from domains with different distributions. The domain is sometimes explicit and other times it is determined by the model [29]. In contrast, this project is focused on explicit hierarchical labels from the same distribution. The objective being to mimic learning different aspects of a single task with left and right hemispheres (2 separate models).
3 Models
We compared a bilateral model to several baselines. Each model consisted of one or more feature extractor, representing hemispheres, as well as one or two classifier heads, one for each label type (fine or coarse). The hemispheres were implemented with two different CNN architectures to test if the effects generalise across architectures: ResNet [35] and VGG [36]. We chose VGG and ResNet because they are common, well understood and amongst the best performing on many vision tasks.
@@ 3 Models
The heads consist of a single fully-connected layer, without bias or non-linearity and with output dimensions the same size as the number of classes (10 for coarse and 100 for fine).
The models are depicted in Figure 1 and described below. Details on training and evaluation in Section 4.1.2.
Figure 1: The range of models. Feature extractors are either ResNet-9 or VGG-11. The classifier heads are single layer fully connected networks, with output dimensions the same size as the number of classes. For brevity, if not specified, the network is unspecialised.
3.1 Bilateral network - specialised
Assembling and training the network consisted of two phases. In Phase 1, individual hemispheres were trained to be specialised feature extractors using supervised learning with a single classifier head, resulting in ‘Unilateral - specialised’, Figure 1a. The left was specialised for specific classes using fine labels, and the right for general classes with coarse labels. In Phase 2, the hemispheres (serving as feature extractors) were brought together into a broader architecture, with two initialised classifier heads, one for fine and coarse labels, resulting in ‘Bilateral - specialised’, Figure 1b, . The weights of the individual hemispheres were frozen and the heads were trained to use the features for classifying fine and coarse labels. A single loss was calculated by adding the loss from each head.
3.2 Baseline models
Unilateral networks with specialisation
We compared the bilateral network to the individual specialised hemispheres alone, on the tasks for which they were specialised, referred to as Unilateral-specialised, Figure 1a. Specialisation was achieved using supervised training with selective labels as described in Section 3.1.
Unilateral network without specialisation
To investigate specialisation, we compared Unilateral-specialised to a single hemisphere without specialisation, referred to as Unilateral-unspecialised, Figure 1d. The single hemisphere was trained on both types of labels simultaneously, by using two classifier heads and adding the loss. This baseline also helps to understand the effect of complementary loss functions; on a single hemisphere compared to the bilateral network.
Bilateral network without specialisation
@@ 3.2 Baseline models
Ensembles
To understand the differences between differential specialisation and conventional ensembling, we compared to a 2-model ensemble, Figure 1e. We chose 2 models, so that the ensemble has approximately the same number of trainable parameters as the bilateral network. We used two unilateral models without specialisation, trained from different seeds. The output of the ensemble was the average from the heads: for fine classification, the average of the fine heads, and for coarse, the average of the coarse heads. We trained five models with different seeds, and then constructed five 2-model ensembles, each with a unique pair of models selected from the pool of five.
4 Experiments
4.1 Experimental setup
4.1.1 Dataset
We used the CIFAR-100 dataset [30], as it includes hierarchical labels that denote fine (specific) and coarse (general) classes. The dataset is split into 50,000 training and 10,000 test samples. Images were resized to 32x32 and we used two widely adopted data augmentation techniques, a random crop and a random horizontal flip (using the inbuilt PyTorch functions) to increase the diversity of the training set.
4.1.2 Training and evaluation
The models were trained with supervised learning using a cross-entropy loss function. In the case of dual heads (one with each label type, see Section 3), the loss from each head was added to create a single loss value. Weights were initialised with PyTorch defaults, from a random uniform distribution $U(-\sqrt{k},\sqrt{k})$: for linear layers $k=1/in\_features$, and for the convolutional layers $k=\frac{1}{C_{in}*\Pi_{i=0}^{1}kernel\_size[i]}$. We used weight decay of 1.0e-5, dropout of 0.6 on the classifier heads to regularise the models, Adam [39] for optimisation, with a learning rate of 1.0e-4 and a mini-batch size of 256.
@@ 4.2 Visualising network operation
Gradient camera (Grad-Cam) visualisation
@@
Feature analysis using cosine similarity
@@
We first grouped the image samples into random pairs with the same class label. We then plotted a bivariate distribution of cosine similarity between the pairs, one for the input to the heads, denoted ‘concatenated’ and the other for the average of the network head outputs, denoted ‘bilateral’. Similarity is plotted along the following dimensions: left hemisphere, right hemisphere and a third dimension, either concatenated or bilateral. Additionally, univariate marginal distributions were plotted for each hemisphere.
@@ 5.2 Visualising network operation
The Grad-Cam visualisations are shown in Figures 4 to 7 for the scenarios described in Section 4.2. In general, the features appear to be more local in the left hemisphere and more global in the right. The heads blend different aspects of features from the left and right, for the task. The effect is that in many cases, features trained for fine classes are helpful for coarse classes and vice versa, and a correct classification can be achieved even if both hemispheres are individually incorrect.
@@ 6 Discussion
Surprisingly, unspecialised hemispheres (Unilateral) performed as well or better than specialised hemispheres (Unilateral-specialised) in most cases. Training with two objective functions (fine and coarse) led to better performance on both labels – whether that is in a bilateral or unilateral architecture.
Ultimately, two hemispheres arranged into a conventional ensemble achieved the best results. A key difference with Bilateral-specialised, is that the component models were more effective to begin with. Another difference is the way in which the model outputs are combined – averaging of the classification vs concatenation of the feature vector. Future work could explore this in more detail.
Both the specialised bilateral network and the ensemble were able to boost performance of individual hemispheres. However, they use different mechanisms.
The novel finding of the paper is that a bilateral architecture can exploit specialised hemispheres, and the remainder of the paper examines this principle.
Computational cost
Compared to a conventional single network, the bilateral architecture has approximately double the number of parameters and computational cost. The performance boost between unilateral and specialised bilateral models may appear modest for the additional cost. However, incremental gains are difficult to achieve at high levels of performance, making the  2.5-5% improvement (across fine, coarse and hemisphere architectures) consequential. This is the first study to demonstrate the bilateral principle, and like other new architectures, greater improvements can be expected in the future.
How does the bilateral architecture work?
The Grad-Cam images reveal that the left hemisphere extracts more localised features than the right. Different learning objectives enable them to capture different aspects of the environment. Collectively the set of features is greater than one network with one objective. Interestingly, even though the left is explicitly trained on fine labels, the features that it extracts are helpful for coarse classes. The inverse is true of the right.
@@
In summary, specialisation creates a higher diversity of features. The network heads implement a type of weighted attention to left and right hemispheres selectively in a task dependent manner, improving overall class prediction.
Relation to biology
A limitation of our model, in terms of biological accuracy, was that inter-hemispheric connectivity was achieved at the outputs of the hemispheres. In contrast, biological hemispheres are interconnected throughout their hierarchies [45]. Nevertheless, the heads may serve a similar purpose, albeit in a cruder way. Inter-hemispheric connectivity comprises a complex combination of inhibitory and excitatory projections. The hemisphere that is better able to represent the input is likely to have stronger activation and thus inhibit the other hemisphere [1, 2]. Like the heads, this too is a type of selectivity. Building a model with interconnected hemispheres throughout the hierarchy is a topic for future research.
The central finding, that left and right hemispheres extracted local and non-local features, to improve performance on a given task, provides one plausible suggestion (and testable hypothesis) for the prevalence of bilateral brains in nature. A possible approach to investigate this idea is with transcranial magnetic stimulation (TMS) to selectively impair one hemisphere at a time on controlled tasks [46].
@@ 7 Conclusion
Inspired by our bilateral brains, we built a bilateral architecture with left and right neural networks using two CNN backbones and compared them to baselines that captured distinct characteristics of the proposed bilateral network. We used a classification task with hierarchical classes that captured an observed characteristic of biological hemispheres, that the left is more specialised for fine classes and the right for coarse classes. The hemispheres were trained to specialise on fine and coarse classes accordingly. They extracted specialised features and through a type of ‘weighted attention’ by a simple fully connected layer, outperformed various (but not all) baselines on classification of both fine and coarse classes. The specialised representations had benefits above the explicit objective of their individual hemispheres. The results demonstrate that small procedural changes to training can achieve specialisation, and that specialisation can be complementary and beneficial for certain tasks. The operation of the artificial network provides ideas for the study of neurobiological hemispheric specialisation. Simultaneously, this work shows that neuroscientific principles can provide inductive biases for novel ML architectures. Currently, in the field of AI/ML where scaling of existing architectures is achieving great success, it’s interesting to look at new principles on smaller architecture that could then be scaled.
Acknowledgements

### v9

# https://cerenaut.ai

@@ Abstract
The brains of all bilaterally symmetric animals on Earth are divided into left and right hemispheres. The anatomy and functionality of the hemispheres have a large degree of overlap, but there are asymmetries, and they specialise in possesses different attributes. Other authors have used computational models to mimic hemispheric asymmetries with a focus on reproducing human data on semantic and visual processing tasks. We took a different approach and aimed to understand how dual hemispheres in a bilateral architecture interact to perform well in a given task. We propose a bilateral artificial neural network that imitates lateralisation observed in nature: that the left hemisphere specialises in local features and the right in global features. We used different training objectives to achieve the desired specialisation and tested it on an image classification task with two different CNN backbones: ResNet and VGG. Our analysis found that the hemispheres represent complementary features that are exploited by a network head that implements a type of weighted attention. The bilateral architecture outperformed a range of baselines of similar representational capacity that do not exploit differential specialisation, with the exception of a conventional ensemble of unilateral networks trained on dual training objectives for local and global features. The results demonstrate the efficacy of bilateralism, contribute to the discussion of bilateralism in biological brains, and the principle may serve as an inductive bias for new AI systems.
Keywords:
Hemispheric specialisation Hemispheric asymmetry Brain-inspired architecture Cognitive architectures Inductive bias
@@ 1 Introduction
Division of the brain into left and right hemispheres is a remarkably conserved feature of brains across species, suggesting the importance of the bilateral architecture for intelligence. The anatomy and physiology of the hemispheres have a large degree of overlap, but they are asymmetric and specialise to possess different attributes. A likely cause of specialisation is the hemispheric asymmetry, which includes differences in neuron density [xweems_hemispheric_2004, 10], distribution of neurotransmitters [30], firing thresholds [30] and intrahemispheric connectivity patterns [32, 26].
One of the emergent functional differences is a specialisation for local (left hemisphere) vs global (right hemisphere) features for visual and auditory processing [13]. In a classic experimental setup [23], a large letter is made up of smaller letters. Subjects are faster at detecting the small letters when presented to the left hemisphere, and the large letters when presented to the right.
Computational models that mimic neuroanatomical and physiological asymmetries are typically subjected to standard tasks and compared to human behavioural data. They provide evidence that asymmetries can cause observed lateralisation of activity for specific tasks, e.g., [29, 22, 32, 37]. The central question is usually to understand the cause of lateralisation/specialisation. However, none have been applied to object recognition, and none have tried to understand how differences can be exploited to improve performance on a given task.
The aim of this project is to investigate the functional benefit of complementary hemispheres for a given task. Our objective is twofold. First, improve performance in a vision task using a bilateral architecture. Second, better understand the functional benefit of a bi-hemispheric architecture and understand how bilateral principles could be exploited more generally for AI/ML.
We focused on the left hemisphere’s specialisation in local features and the right’s in global features. To investigate this, we chose a hierarchical image classification task where each sample belongs to both a fine and a coarse class, e.g. coarse: sea creature, fine: penguin, seal, shark. In this setup, fine labels represent local features while coarse labels represent global features.
Our approach is to construct a bilateral network and induce specialisation in each hemisphere directly, rather than to test if it emerges as a result of architectural asymmetry. We modelled the hemispheres with left and right CNNs and induced specialisation using supervised training with different objectives, using either coarse or fine labels. The bilateral network was compared with several baselines that do not have specialisation but do have the same number of learnable parameters. The differences were analysed to explore the effects of bilateral specialisation. Experiments were repeated with two types of CNN hemispheres, VGG and ResNet variants, to test generalisation over different backbones.
@@ 2 Related work
Specialisation for local and global visual features has been attributed to differential sensitivity to spatial frequencies and explored with computational models for face perception [7, 14, 37] and a line bisection task [34, 21]. However, other studies do not support hemispheric specialisation for particular frequency ranges, and Hsiao et al. [13] showed that this specialisation may emerge as a result of differences in neuron connectivity.
Beaulieu et al. (2020) [3] created a dual-stream architecture called Neuromodulator Meta-learner, where one network learns to modulate the other, to enhance continual learning. Bakhtiari et al. (2021) [2] created a network with parallel pathways to reproduce the functionality of dorsal (‘where’) and ventral (‘what’) pathways in an ANN trained with a single loss function. Li et al. (2021) [19] discussed specialisation of a branched neural network.
In the most related study, Mayan et al. (2021) [20] tackled the same classification task with fine and coarse (hierarchical) labels. A single network with two hemispheres was trained with supervised learning and a single loss function. They used different hyperparameters in each hemisphere, with analogies to biological parameters, to encourage specialisation; which was achieved but the bilateral network did not have an advantage over baselines.
@@
In multi-domain learning (MDL), multiple models are trained to specialise in a particular domain [12]. Two of the main approaches include a) modelling domain-specific and domain-general features [15] and b) explicitly modelling the relationship between domains. MDL is focused on classes from domains with different distributions. The domain is sometimes explicit and other times it is determined by the model [6]. In contrast, this project is focused on explicit hierarchical labels from the same distribution. The objective being to mimic learning different aspects of a single task with left and right hemispheres.
3 Methods
3.1 Dataset
We used the CIFAR-100 dataset [17], as it includes hierarchical labels that denote fine (local) and coarse (global) classes. There are 100 fine and 20 coarse classes. The dataset is split into 50,000 training and 10,000 test samples. Images were resized to 32x32 and we used two widely adopted data augmentation techniques, a random crop and a random horizontal flip (using inbuilt PyTorch functions) to increase diversity of the training set.
3.2 Models and specialisation approach
We compared a bilateral model to several baselines, shown in Figure 1 and described below. Details on training and evaluation in Section 3.3. Each model consisted of one or more feature extractor, representing hemispheres, as well as one or two classifier heads, one for each label type (fine or coarse). The hemispheres were implemented with two different CNN architectures to test if the effects generalise across architectures: ResNet [11] and VGG [35]. We chose VGG and ResNet because they are common, well understood and amongst the best performing on many vision tasks.
Figure 1: Bilateral architecture and baselines. Feature extractors are either ResNet-9 or VGG-11. The classifier heads are single layer fully connected networks, with output dimensions the same size as the number of classes. For brevity, if not specified, the network is unspecialised.
3.2.1 Common elements – backbones and heads
@@
The heads consist of a single fully-connected layer, without bias or non-linearity and with dimensions the same size as the number of classes (10 for coarse and 100 for fine).
3.2.2 Bilateral network with specialisation
Assembling and training the network consisted of two phases. In Phase 1, individual hemispheres were trained to be specialised feature extractors using supervised learning with a single classifier head, resulting in ‘Unilateral - specialised’, Figure 1a. The left was specialised for local features using fine labels, and the right was specialised for global features with coarse labels. In Phase 2, the hemispheres (serving as feature extractors) were brought together into a broader architecture, with two initialised classifier heads, one for fine and coarse labels, resulting in ‘Bilateral - specialised’, Figure 1b, . The weights of the individual hemispheres were frozen and the heads were trained to use the features for classifying fine and coarse labels. A single loss was calculated by adding the loss from each head.
3.2.3 Unilateral network with specialisation
We compared the bilateral network with the individual specialised hemispheres alone, on the tasks for which they were specialised, referred to as Unilateral-specialised, Figure 1a.
3.2.4 Unilateral network without specialisation
To investigate specialisation, we compared Unilateral-specialised to a single hemisphere without specialisation, referred to as Unilateral-unspecialised, Figure 1d. The single hemisphere was trained on both types of labels simultaneously, using two classifier heads and adding the loss. This baseline also helps to understand the effect of complementary loss functions; on a single hemisphere compared to the bilateral network.
3.2.5 Bilateral network without specialisation
@@
3.2.6 Ensembles
To understand the differences between differential specialisation and conventional ensembling, we compared to a 2-model ensemble, Figure 1e. We chose 2 models, so that the ensemble has approximately the same number of trainable parameters as the bilateral network. We used two unilateral models without specialisation, trained from different seeds. The output of the ensemble was the average of the heads: for fine classification, the average of the fine heads, and for coarse, the average of the coarse heads. We trained five models with different seeds, and then constructed five 2-model ensembles, each with a unique pair of models selected from the pool of five.
3.3 Training procedure
The models were trained with supervised learning using a cross-entropy loss function. In the case of dual heads, the loss from each head was added to create a single loss value. Weights were initialised with PyTorch defaults, from a random uniform distribution $U(-\sqrt{k},\sqrt{k})$: for linear layers $k=1/in\_features$, and for convolutional layers $k=\frac{1}{C_{in}*\Pi_{i=0}^{1}kernel\_size[i]}$. We used weight decay of 1.0e-5, dropout of 0.6 on the classifier heads to regularise the models, Adam optimiser [16], a learning rate of 1.0e-4 and mini-batch size of 256. These hyperparameters were empirically optimised with manual coordinate descent.
@@ 3.4 Visualising network operation
3.4.1 Gradient camera (Grad-Cam) visualisation
@@
3.4.2 Feature analysis using cosine similarity
@@
We first grouped the image samples into random pairs with the same class label. We then plotted a bivariate distribution of cosine similarity between the pairs, one for the input to the heads, denoted ‘concatenated’ and the other for the average of the network head outputs, denoted ‘bilateral’. Similarity is plotted along the following dimensions: left hemisphere, right hemisphere and a third dimension, either concatenated or bilateral. In addition, univariate marginal distributions were plotted for each hemisphere.
3.5 Experimental setup
@@ 4.2 Visualising network operation
Grad-Cam visualisations are shown in Figures 4 to 7 for scenarios described in Section 3.4. In general, the features appear to be more local in the left hemisphere and more global in the right. The heads blend different aspects of features from the left and right, for the task. The effect is that in many cases, features trained for fine classes are helpful for coarse classes and vice versa, and a correct classification can be achieved even if both hemispheres are individually incorrect.
@@ 5 Discussion
Surprisingly, unspecialised hemispheres (Unilateral) performed as well as or better than specialised hemispheres (Unilateral-specialised) in most cases. Recall that unilateral models are trained with two objective functions (fine and coarse). So using two training objectives led to better performance on both label types, whether that is in a bilateral or unilateral architecture. Using two explicit objectives could be seen as a form of specialisation, even when used simultaneously and in a single model.
Ultimately, two hemispheres arranged in a conventional ensemble achieved the best results. A key difference with Bilateral-specialised is that the component models of the ensemble were more effective to begin with; benefiting from two training objectives. Another difference is the way that the model outputs are combined – averaging of the classification vs concatenation of the feature vector. Future work could explore this in more detail.
Both the specialised bilateral network and the ensemble were able to boost performance of individual hemispheres. However, they use different mechanisms. The novel finding of the paper is that a bilateral architecture can exploit specialised hemispheres, and the remainder of the paper examines this principle.
5.0.1 Computational cost
The performance boost of bilateral models over a single hemisphere may appear modest relative to the additional computational cost, which could limit their practical applications. However, incremental gains are difficult to achieve at high levels of performance, making the  2.5-5% improvement (across fine, coarse and hemisphere architectures) consequential. Like other new architectures, improvements can be expected in the future.
5.0.2 How does the bilateral architecture work?
The Grad-Cam images reveal that the left hemisphere extracts more localised features than the right. Different learning objectives enable them to capture different aspects of the environment. Collectively, the set of features is greater than one network with one objective. Interestingly, even though the left is explicitly trained on fine labels, the features it extracts are helpful for coarse classes. The inverse is also true.
@@ 5.0.2 How does the bilateral architecture work?
In summary, specialisation creates a higher diversity of features. The network heads implement a type of weighted attention to left and right hemispheres selectively in a task-dependent manner, improving overall class prediction.
5.0.3 Relation to biology
A limitation in terms of biological accuracy was that interhemispheric connectivity was achieved at the outputs of the hemispheres. In contrast, biological hemispheres are interconnected throughout their hierarchies [4]. However, heads may serve a similar purpose, albeit in a cruder way. Interhemispheric connectivity comprises a complex combination of inhibitory and excitatory projections. The hemisphere that is better able to represent the input is likely to have stronger activation and thus inhibit the other hemisphere [38]. Like the heads, this too is a type of selectivity. Interconnected hemispheres throughout the hierarchy is a topic for future work.
The main finding, that left and right hemispheres extracted local and non-local features to improve performance on a given task, provides one plausible reason (and testable hypothesis) for the prevalence of bilateral brains in nature. A possible approach to investigate this is with transcranial magnetic stimulation (TMS) to selectively impair one hemisphere at a time on controlled tasks [27].
@@ 6 Conclusion
Inspired by our bilateral brains, we built a bilateral architecture with left and right neural networks using two CNN backbones and compared them to baselines that captured distinct characteristics of the proposed bilateral network. We used a classification task with hierarchical classes that captured an observed characteristic of biological hemispheres, that the left is more specialised for fine classes and the right for coarse classes. The hemispheres were trained to specialise on fine and coarse classes accordingly. They extracted specialised features and through a type of ‘weighted attention’ by a simple fully connected layer, outperformed various (but not all) baselines on classification of both fine and coarse classes. The specialised representations had benefits above the explicit objective of their individual hemispheres. The results demonstrate that small procedural changes to training can achieve specialisation, and that specialisation can be complementary and beneficial for certain tasks. The operation of the artificial network provides ideas for the study of neurobiological hemispheric specialisation. Simultaneously, this work shows that neuroscientific principles can provide inductive biases for novel ML architectures. Currently, in AI/ML where scaling of existing architectures is achieving great success, it’s interesting to look at new principles on smaller architecture that could then be scaled.
{credits}
6.0.1 Acknowledgements